# Context-Guided Spatial Feature Reconstruction for Efficient Semantic Segmentation

### v1

@@ Abstract
Semantic segmentation is an important task for many applications but it is still quite challenging to achieve advanced performance with limited computational costs. In this paper, we present CGRSeg, an efficient yet competitive segmentation framework based on context-guided spatial feature reconstruction. A Rectangular Self-Calibration Module is carefully designed for spatial feature reconstruction and pyramid context extraction. It captures the global context in both horizontal and vertical directions and gets the axial global context to explicitly model rectangular key areas. A shape self-calibration function is designed to make the key areas more close to the foreground object. Besides, a lightweight Dynamic Prototype Guided head is proposed to improve the classification of foreground objects by explicit class embedding. Our CGRSeg is extensively evaluated on ADE20K, COCO-Stuff, and Pascal Context benchmarks, and achieves state-of-the-art semantic performance. Specifically, it achieves $43.6\%$ mIoU on ADE20K with only $4.0$ GFLOPs, which is $0.9\%$ and $2.5\%$ mIoU better than SeaFormer and SegNeXt but with about $38.0\%$ fewer GFLOPs. Code is available at https://github.com/nizhenliang/CGRSeg.
@@ 1 Introduction
Semantic segmentation is a foundational computer vision task that aims to assign semantic labels to each pixel in an image. In recent years, deep learning methods have achieved remarkable results in semantic segmentation, thanks to the availability of large-scale datasets, powerful computing resources, and advanced network architectures. Early studies for semantic segmentation are mainly based on convolution neural networks (CNNs), such as FCN [23], PSPNet [36], and DeepLab [4, 6, 5, 21]. More recently, a series of methods for semantic segmentation have been proposed based on vision transformer and achieve competitive performance [37, 28, 33, 25]. Deep learning based methods have significantly improved the accuracy of semantic segmentation. However, it is still a quite challenging topic for how to achieve advanced segmentation performance with limited computing resources.
@@
Efficient and lightweight semantic segmentation is an important topic and has attracted great research interest. Yu et al. [29] proposed BiSeNet, which is a two-pathway architecture to capture local and global information and achieves a good balance between speed and segmentation performance. SegFormer [28] was built upon a hierarchically structured Transformer encoder and a lightweight multilayer perception (MLP) decoder to achieve high precision and running speeds. Wan et al.  [25] introduced an efficient squeeze-enhanced Axial Transformer (SeaFormer) to reduce the computational complexity of axial attention. Recently, Guo et al.  [14] proposed SegNeXt to demonstrate that an encoder with simple and cheap convolutions can still achieve competitive performance.
@@
Nevertheless, how to make the lightweight model have higher segmentation accuracy is still to be explored. Due to the limited feature representation capability, it is difficult for lightweight models to model the boundaries and categories of foreground objects, which can lead to problems such as inaccurate boundary segmentation and misclassification. To address these problems, we carefully design the Rectangular Self-Calibration Module (RCM) to improve the position modeling of foreground objects. Besides, a Dynamic Prototype Guided (DPG) head is proposed to embed class information to improve the class discrimination of foreground objects. Moreover, to further introduce the pyramid context to improve the feature representation, we design a context-guided spatial feature reconstruction network (CGRSeg), which consists of pyramid context extraction, spatial feature reconstruction, and a lightweight head.
@@
The Rectangular Self-Calibration Module is proposed to improve the ability to locate foreground objects and extract pyramid context. The rectangular self-calibration attention is a core component of RCM. It adopts horizontal pooling and vertical pooling to capture the axial global context and generate two axis vectors. The two axis vectors are added together to model a rectangular attention region. A shape self-calibration function is designed to adjust the shape of rectangular attention and bring it closer to the foreground features, which consists of large-kernel strip convolution. Besides, a fusion function is designed to fuse attention features and increase local detail. In this way, RCM can make the model focus more on the foreground for spatial feature reconstruction and can capture axial global context for pyramid context extraction.
@@
The Dynamic Prototype Guided head is proposed to improve the classification of the foreground objects via explicit class embedding. Specifically, the features are first projected into the class feature space. And, the feature in class space and the feature in pixel space are multiplied to obtain the dynamic prototype. The dynamic prototype can reflect the feature distribution of all classes on each image. To further embed class information, the dynamic prototype is compressed to obtain the class embedding vector. Finally, the class embedding vector is projected into the pixel feature space, and the pixel features are weighted to enhance the discrimination between different classes. In this way, the DPG head can better distinguish features of different classes and improve classification performance.
@@
The proposed CGRSeg achieves state-of-the-art segmentation performance on ADE20K, COCO Stuff, and Pascal Context with fewer FLOPs and higher Throughputs. The performance on ADE20K is shown in Figure 1. Specifically, our method achieves $43.6\%$ mIoU on ADE20K with only $4.0$ GFLOPs, which is $0.9\%$ and $2.5\%$ mIoU better than SeaFormer [28] and SegNeXt [14] but with about $38.0\%$ fewer GFLOPs.
@@ 2 Related Work
Efficient Semantic Segmentation Segmentation is one of the basic tasks in computer vision. The lightweight semantic segmentation methods based on CNN include BiSeNet series [29], STDCNet [11], DFANet [18], SegNeXt [14] and so on. BiSeNet [29] first proposed a two-pathway architecture that uses a spatial path and a context path to capture both local and global information. STDCNet [11] designed a STDC module to extract scalable receptive field and multi-scale information and it uses boundary supervision to supplement spatial details. SegNeXt [14] evoked spatial attention via multi-scale convolutional features. It indicated that an encoder with simple and cheap convolutions can still perform better than vision transformers. In recent years, some Transformer-based methods have been proposed, such as SegFormer [25], TopFormer [35], and SeaFormer [25]. TopFormer [35] proposed a lightweight segmentation network based on CNN and transformer. It takes Tokens from various scales as input to produce scale-aware semantic features, which are then injected into the corresponding tokens to augment the representation. SeaFormer [25] designed a squeeze-enhanced axial Transformer in the encoder to improve feature representation and proposed a new two-path structure. It can achieve high precision with small flops. Different from these methods, our CGRSeg employed pyramid context-guided spatial feature reconstruction to improve the ability to model foreground objects, which is more efficient.
@@
Efficient Context Extraction Modules In previous works, attention is often used for context extraction, such as coordinate attention and GatherExcite. Coordinate attention [15] captured the axial global context through horizontal and vertical pooling and extracted coordinate relationships to enhance positional awareness. GatherExcite [16] efficiently aggregated feature responses from a large spatial extent, and redistributed the pooled information to local features. The proposed RCM uses addition to model the critical region, and designs shape self-calibration to calibrate the rectangle attention, which can make the model pay more attention to the foreground features. These designs allow RCM to outperform existing attention. With the development of transformers, some convolutional networks [22, 30] adopted transformer-like structures to improve feature representation. ConvNeXt [22] first proposed the structure of large kernel depth-wise convolution combined with BN and MLP, which can significantly improve the feature representation of traditional convolutional neural networks. InceptionNeXt [30] proposed the MetaNeXt architecture and combined the Inception block with the MetaNeXt architecture to significantly improve model performance. These works show that the MetaNeXt structure has powerful context extraction ability.
@@
Segmentation Head Segmentation heads are usually designed to improve the performance of segmentation networks. Some segmentation heads use multi-scale features to achieve advanced performance, such as PSPNet [36], DeepLab series [4, 6, 5, 21]. PSPNet [36] designed pyramid pooling model to capture multi-scale information. Some segmentation heads, such as OCRHead [31] and DAHead [12], captured long-range dependencies to improve performance. Hamburger Head [13] is a powerful global context module based on matrix decomposition, surpassing various attention modules on semantic segmentation. However, the existing segmentation heads are not suitable for efficient semantic segmentation because of their high computational complexity.
@@ 3.1 CGRSeg
Previous work [35, 25, 12, 7] has shown that foreground object modeling and pyramid context extraction are critical for segmentation. To this end, we design a context-guided spatial feature reconstruction network (CGRSeg). As is shown in Figure 2, CGRSeg consists of three key parts: pyramid context extraction, spatial feature reconstruction, and a lightweight head. To make the model focus on the foreground features, a Rectangular Self-Calibration Module (RCM) is proposed. Besides, it captures the axial global context for pyramid context extraction. A Dynamic Prototype Guided (DPG) head is proposed to improve the classification of the foreground objects via explicit class embedding.
Figure 2: The overall architecture of CGRSeg. The Rectangular Self-Calibration Module (RCM) is designed for spatial feature reconstruction and pyramid context extraction. The rectangular self-calibration attention (RCA) explicitly models the rectangular region and calibrates the attention shape. The Dynamic Prototype Guided (DPG) head is proposed to improve the classification of the foreground objects via explicit class embedding.
Pyramid Context Extraction The RCM is applied to extract pyramid context, which applies horizontal and vertical pooling to capture axial context in two directions. Besides, RCM employs MLP to further enhance feature representation. As shown in Figure 2, a stepwise downsampling encoder is used in the proposed model. Encoder produces features of different scales features as $[\mathbf{F}_{1}$, $\mathbf{F}_{2}$, $\mathbf{F}_{3}$, $\mathbf{F}_{4}]$, with the resolution of $[\frac{H}{4}\times\frac{W}{4}$,$\frac{H}{8}\times\frac{W}{8}$,$\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{W}{32}]$, respectively. To ensure the efficiency of the whole network, the largest-scale feature, i.e., $\mathbf{F}_{1}$, is dropped by the decoder. Then, the lower-scale features $\mathbf{F}_{2}$, $\mathbf{F}_{3}$ and $\mathbf{F_{4}}$ are down-sampled to $\frac{H}{64}\times\frac{W}{64}$ size by the average pooling, and they are concatenated together to generate the pyramid feature $\mathbf{F}_{5}$. $\mathbf{F}_{5}$ is fed into multiple stacked RCMs for pyramid feature interaction and extract scale-aware semantic features. Finally, the features are split and upsampled to the original scale after the pyramid feature extraction. This process can be formulated as:
@@
where $AP(\mathbf{F},x)$ represents the operation of average pooling that down-samples the feature $\mathbf{F}$ by the factor $x$. $P$ is the feature with pyramid context.
Spatial Feature Reconstruction To make the decoder features more focused on the foreground, RCM is used to reconstruct the spatial features. Specifically, the CGRSeg fuses the low-level spatial features from the encoder with the high-level features of the corresponding scale in the decoder. The fused features are reconstructed by using RCM. RCM captures the axial global context to model the rectangular critical region, and then uses the shape self-calibration function to adjust the attention region to the foreground. Moreover, pyramid features are applied to guide the spatial feature reconstruction to make the reconstructed features feel multi-scale information.
@@ 3.2 Rectangular Self-Calibration Module
The Rectangular Self-Calibration Module (RCM) is designed to make the model focus on the foreground features. It can also capture the axial global context for pyramid context extraction. The module consists of rectangular self-calibration attention, batch normalization, and MLP.
Figure 3: The shape change of the highlighted region is caused by the rectangular self-calibration attention. By training to adjust the weights of the two strip convolutions, the attention region is calibrated closer to the foreground object.
The rectangular self-calibration attention adopts horizontal pooling and vertical pooling to capture the axial global context in two directions and uses broadcast addition to model the rectangle region of interest. Then, a shape self-calibration function is designed to calibrate the region of interest, which can make the region of interest closer to the foreground object. Here, two large-kernel strip convolutions are used to calibrate the attention map in the horizontal and vertical directions decoupledly. First, we use the horizontal strip convolution to calibrate the shape in the horizontal direction, which adjusts each row of elements to make the horizontal shape closer to the foreground object. Then, the features are normalized by BN and the non-linearity is added by ReLU. The shape is also calibrated by using the vertical strip convolution in the vertical direction. In this way, the convolution in two directions can be decoupled, which can adapt to any shape.
The weights of the strip convolutions are learnable. As shown in Figure 3, the shape of the highlighted area is changed by two strip convolutions with different weights. By training, this function can learn the appropriate weights to adjust the rectangular area to the foreground object. Furthermore, in Figure 2, the visualization results show that the rectangular self-calibration attention does effectively model the rectangular key region and adjusts the features to focus more on the foreground by the shape self-calibration function. The shape self-calibration function can be formulated as follows:
$\xi_{C}(\overline{y})=\delta(\psi_{k\times 1}(\phi(\psi_{1\times k}(\overline{y}))))$ (2)
where $\psi$ indicates the large kernel strip convolution, and $k$ indicates the kernel length of the strip convolution. $\phi$ indicates the Batch Normalization followed by the ReLU function, and $\delta$ indicates the Sigmoid function.
Besides, a feature fusion function is designed to fuse attention features with input features. Specifically, 3 $\times$ 3 depth-wise convolution is used to further extract local details of the input feature, and the calibrated attention feature is weighted to the feature by Hadamard Product.
$\xi_{F}(x,y)=\psi_{3\times 3}(x)\odot y$ (3)
@@
Recently, ConvNeXt [22] proposed the MetaNeXt structure which significantly improves the performance of convolution in the classification task. Inspired by this idea, we combine rectangular self-calibration attention with the MetaNeXt structure to further improve feature representation. Specifically, batch normalization and MLP are added after the rectangular self-calibration attention to refine features. Finally, the residual connection is employed to further enhance feature reuse. The architecture of the RCM is shown in Figure 2. It also can be described as:
$\rm\mathbf{F}_{out}=\rho(\xi_{F}(x,\xi_{C}(H_{P}(x)\oplus V_{P}(x))))+\mathbf{F}_{in}$ (4)

### v2

@@ Abstract
Semantic segmentation is an important task for numerous applications but it is still quite challenging to achieve advanced performance with limited computational costs. In this paper, we present CGRSeg, an efficient yet competitive segmentation framework based on context-guided spatial feature reconstruction. A Rectangular Self-Calibration Module is carefully designed for spatial feature reconstruction and pyramid context extraction. It captures the axial global context in both horizontal and vertical directions to explicitly model rectangular key areas. A shape self-calibration function is designed to make the key areas closer to foreground objects. Besides, a lightweight Dynamic Prototype Guided head is proposed to improve the classification of foreground objects by explicit class embedding. Our CGRSeg is extensively evaluated on ADE20K, COCO-Stuff, and Pascal Context benchmarks, and achieves state-of-the-art semantic performance. Specifically, it achieves $43.6\%$ mIoU on ADE20K with only $4.0$ GFLOPs, which is $0.9\%$ and $2.5\%$ mIoU better than SeaFormer and SegNeXt but with about $38.0\%$ fewer GFLOPs. Code is available at https://github.com/nizhenliang/CGRSeg.
@@ 1 Introduction
Semantic segmentation is a foundational computer vision task that aims to assign semantic labels to each pixel in an image. In recent years, deep learning methods have achieved remarkable results in semantic segmentation, thanks to the availability of large-scale datasets, powerful computing resources, and advanced network architectures. Early studies for semantic segmentation are mainly based on convolution neural networks (CNNs), such as FCN [27], PSPNet [43], and DeepLab [6, 8, 7, 25]. More recently, a series of methods for semantic segmentation have been proposed based on vision transformer and achieve competitive performance [44, 35, 40, 30]. Deep learning methods have significantly improved the accuracy of semantic segmentation. However, it is still quite a challenging topic for how to achieve advanced segmentation performance with limited computing resources.
@@
Efficient and lightweight semantic segmentation is an important topic and has attracted great research interest. Yu et al. [36] introduced a two-pathway architecture named BiSeNet, which captures both local and global information and achieves a good balance between speed and performance. SegFormer [35] designed a hierarchical Transformer encoder coupled with a lightweight multilayer perception (MLP) decoder to achieve both high precision and speeds. Wan et al.  [30] introduced an efficient squeeze-enhanced axial transformer (SeaFormer) to reduce the computational complexity of the standard transformer. Recently, Guo et al.  [16] proposed SegNeXt to demonstrate that an encoder with simple and cheap convolutions can still achieve competitive performance.
@@
Nevertheless, existing lightweight models, constrained by their limited feature representation capabilities, often encounter difficulties in modeling the boundaries and distinguishing the categories of foreground objects. These limitations lead to inaccuracies in boundary segmentation and misclassification. To address these problems, we carefully design the Rectangular Self-Calibration Module (RCM) to improve the position modeling of foreground objects. Besides, a Dynamic Prototype Guided (DPG) head is introduced to embed class information to enhance the class discrimination of foreground objects. Moreover, to further introduce the pyramid context to improve the feature representation, we design a context-guided spatial feature reconstruction network (CGRSeg) consisting of pyramid context extraction, spatial feature reconstruction, and a lightweight head.
@@
The Rectangular Self-Calibration Module is proposed to improve the ability to locate foreground objects and extract pyramid context. The rectangular self-calibration attention is a core component of RCM. It adopts horizontal pooling and vertical pooling to capture the axial global context and generate two axis vectors. The two axis vectors are added together to model a rectangular attention region. A shape self-calibration function is designed to adjust the shape of rectangular attention and bring it closer to the foreground feature, which consists of large-kernel strip convolution. Besides, a fusion function is designed to fuse attention features and increase local detail. In this way, RCM can make the model focus more on the foreground for spatial feature reconstruction, while also effectively capturing the axial global context for pyramid context extraction.
@@
The Dynamic Prototype Guided head is proposed to improve the classification of the foreground objects via explicit class embedding. Specifically, the features are first projected into the class feature space. Then, the feature in class space and the feature in pixel space are multiplied to obtain the dynamic prototype. The dynamic prototype can reflect the feature distribution of all classes on each image. To further embed class information, the dynamic prototype is compressed to obtain the class embedding vector. Finally, the class embedding vector is projected into the pixel feature space, and the pixel features are weighted to enhance the discrimination between different classes. In this way, the DPG head can effectively enhance class discrimination and improve classification performance.
@@
The proposed CGRSeg achieves state-of-the-art segmentation performance on ADE20K, COCO Stuff, and Pascal Context with fewer FLOPs and higher Throughput. The performance on ADE20K is shown in Fig. 1. Specifically, our method achieves $43.6\%$ mIoU on ADE20K with only $4.0$ GFLOPs, which is $0.9\%$ and $2.5\%$ mIoU better than SeaFormer [35] and SegNeXt [16] but with about $38.0\%$ fewer GFLOPs.
@@ 2 Related Work
Efficient Semantic Segmentation The lightweight semantic segmentation methods based on CNN include BiSeNet series [36], STDCNet [13], DFANet [20], SegNeXt [16] and so on. STDCNet [13] designed an STDC module to extract scalable receptive field and multi-scale information and used boundary supervision to supplement spatial details. SegNeXt [16] captured spatial attention via multi-scale convolutional features, achieving better performance than vision transformers. In recent years, several Transformer-based methods have emerged in this field [24, 30, 42, 35]. TopFormer [42] introduced a hybrid architecture of CNN and Transformer, which took tokens from various scales as input to produce scale-aware semantic features. SeaFormer [30] designed a squeeze-enhanced axial transformer in the encoder to improve feature representation and proposed a new two-path structure. Besides, there are some methods to improve the segmentation performance through contrast learning or other training approaches, such as HSSN [21], and ContrastiveSeg [32]. Different from these methods, our CGRSeg employs pyramid context-guided spatial feature reconstruction to improve the ability to model foreground objects, which is more efficient.
@@
Efficient Context Extraction Modules In previous works, attention is often used for context extraction [17, 18, 4, 33], such as coordinate attention and GatherExcite. Coordinate attention [17] captured the axial global context through horizontal and vertical pooling, which also extracted coordinate relationships to enhance positional awareness. GatherExcite [18] efficiently aggregated contextual information across large neighborhoods to refine the feature representation. The proposed RCM uses addition to model the critical region, and designs shape self-calibration to calibrate the rectangle attention, which can make the model pay more attention to the foreground features. These designs allow RCM to outperform existing attention. With the development of transformers, some convolutional networks [26, 37] adopted transformer-like structures to enhance feature representation. ConvNeXt [26] first proposed the structure of large kernel depth-wise convolution combined with BN and MLP, which can significantly improve the feature representation of traditional convolutional neural networks. InceptionNeXt [37] proposed the MetaNeXt architecture and combined the Inception block with the MetaNeXt architecture, significantly improving the model performance. These works show that the MetaNeXt structure has powerful context extraction ability.
@@
Segmentation Head Segmentation heads are usually designed to improve the performance of segmentation networks. Some segmentation heads use multi-scale features to achieve advanced performance, such as PSPNet [43], DeepLab series [6, 8, 7, 25]. PSPNet [43] designed the pyramid pooling module to capture multi-scale information. Some segmentation heads, such as OCRHead [38] and DAHead [14], captured long-range dependencies to improve performance. Hamburger Head [15] introduced a powerful global context module based on matrix decomposition, surpassing various attention modules on semantic segmentation. There is also some work to improve the model performance by improving the classifier [46]. ProtoSeg [46] proposed a new perspective on semantic segmentation by replacing the pixel-wise classifier with the prototype-based classifier. However, existing segmentation heads are not suitable for efficient segmentation because they require large computational resources and have poorer performance.
@@ 3.1 CGRSeg
Previous work [42, 30, 14, 9] has shown that foreground object modeling and pyramid context extraction are critical for segmentation. To this end, we design a context-guided spatial feature reconstruction network (CGRSeg). As shown in Fig. 2, CGRSeg consists of three key parts: pyramid context extraction, spatial feature reconstruction, and a lightweight head. To make the model focus on the foreground features, a Rectangular Self-Calibration Module (RCM) is proposed. Besides, it captures the axial global context for pyramid context extraction. A Dynamic Prototype Guided (DPG) head is proposed to improve the classification of the foreground objects via explicit class embedding.
Figure 2: The overall architecture of CGRSeg. The Rectangular Self-Calibration Module (RCM) is designed for spatial feature reconstruction and pyramid context extraction. The rectangular self-calibration attention (RCA) explicitly models the rectangular region and calibrates the attention shape. The Dynamic Prototype Guided (DPG) head improves the classification of the foreground objects via explicit class embedding.
Pyramid Context Extraction The RCM is applied to extract pyramid context, which applies horizontal and vertical pooling to capture the axial context in two directions. Besides, RCM employs MLP to further enhance feature representation. As shown in Fig. 2, a stepwise downsampling encoder is used in the proposed model. Encoder produces features of different scales features as $[\mathbf{F}_{1}$, $\mathbf{F}_{2}$, $\mathbf{F}_{3}$, $\mathbf{F}_{4}]$, with the resolution of $[\frac{H}{4}\times\frac{W}{4}$,$\frac{H}{8}\times\frac{W}{8}$,$\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{W}{32}]$, respectively. To ensure the efficiency of the whole network, the largest-scale feature, i.e., $\mathbf{F}_{1}$, is dropped by the decoder. Then, the lower-scale features $\mathbf{F}_{2}$, $\mathbf{F}_{3}$ and $\mathbf{F_{4}}$ are down-sampled to $\frac{H}{64}\times\frac{W}{64}$ size by the average pooling, and they are concatenated together to generate the pyramid feature $\mathbf{F}_{5}$. $\mathbf{F}_{5}$ is fed into multiple stacked RCMs for pyramid feature interaction and extract scale-aware semantic features. Finally, the features are split and upsampled to the original scale after the pyramid feature extraction. This process can be formulated as:
@@
where $AP(\mathbf{F},x)$ represents the average pooling that down-samples the feature $\mathbf{F}$ by the factor $x$. $P$ is the feature with pyramid context.
Spatial Feature Reconstruction To make the decoder features more focused on the foreground, RCM is employed to reconstruct the spatial features. Specifically, the CGRSeg fuses the low-level spatial features from the encoder with the high-level features of the corresponding scale in the decoder. The fused features are reconstructed by using RCM. RCM captures the axial global context to model the rectangular critical region and then uses the shape self-calibration function to adjust the attention region to the foreground. Moreover, pyramid features are applied to guide the spatial feature reconstruction to make the reconstructed features feel multi-scale information.
@@ 3.2 Rectangular Self-Calibration Module
The Rectangular Self-Calibration Module (RCM) is designed to make the model focus on the foreground. It can also capture the axial global context for pyramid context extraction. The module consists of rectangular self-calibration attention, batch normalization, and MLP.
Figure 3: The shape change of the highlighted region is caused by the rectangular self-calibration attention. By optimizing the weights of the two strip convolutions during training, the attention region is calibrated closer to the foreground object.
The rectangular self-calibration attention (RCA) adopts horizontal pooling and vertical pooling to capture the axial global context in two directions, generating two distinct axis vectors. By applying broadcast addition to these two vectors, RCA can effectively model the rectangle region of interest. Then, a shape self-calibration function is designed to calibrate the region of interest, which can make the region of interest closer to the foreground object. Here, two large-kernel strip convolutions are utilized to calibrate the attention map in the horizontal and vertical directions decoupledly. First, we employ the horizontal strip convolution to calibrate the shape in the horizontal direction, which adjusts each row of elements to make the horizontal shape closer to the foreground object. Then, the features are normalized by BN, and the non-linearity is added by ReLU. Subsequently, the shape is also calibrated by using the vertical strip convolution in the vertical direction. In this way, the convolution in two directions can be decoupled, which can adapt to any shape.
The weights of the strip convolutions are learnable. As shown in Fig. 3, the shape of the highlighted area is changed by two strip convolutions with different weights. By training, this function can learn the appropriate weights to adjust the rectangular area to the foreground object. Furthermore, in Fig. 2, the visualization results show that the rectangular self-calibration attention does effectively model the rectangular key region and adjusts the features to focus more on the foreground by the shape self-calibration function. The shape self-calibration function can be formulated as follows:
$\xi_{C}(\overline{y})=\delta(\psi_{k\times 1}(\phi(\psi_{1\times k}(\overline{y})))),$ (2)
where $\psi$ indicates the large-kernel strip convolution, and $k$ indicates the kernel size of the strip convolution. $\phi$ indicates the Batch Normalization followed by the ReLU function, and $\delta$ indicates the Sigmoid function.
Besides, a feature fusion function is designed to fuse attention features with input features. Specifically, 3 $\times$ 3 depth-wise convolution is used to further extract local details of the input feature. The calibrated attention feature is weighted to the refined input feature by Hadamard Product.
$\xi_{F}(x,y)=\psi_{3\times 3}(x)\odot y,$ (3)
@@
Recently, ConvNeXt [26] proposed the MetaNeXt structure which significantly improves the performance of the convolutional neural network in the classification task. Inspired by this idea, we combine rectangular self-calibration attention with the MetaNeXt structure to further improve feature representation. Specifically, batch normalization and MLP are added after the rectangular self-calibration attention to refine features. Finally, the residual connection is employed to further enhance feature reuse. The architecture of the RCM is shown in Fig. 2. It also can be described as:
$\rm\mathbf{F}_{out}=\rho(\xi_{F}(x,\xi_{C}(H_{P}(x)\oplus V_{P}(x))))+x,$ (4)